Please answer this problem using systems of linear equations in three variables.
The sum of the length, width and the height of a rectangular box is 17 cm. Twice the width exceeds the height by by 1 and the sum of the length and the width is shorter than the height by 1 cm. Find the dimensions of the box.

Answer by htmentor(1343)

l+w+h=17 (1)
2w=h+1 (2)
l+w=h-1 (3)
Substitute the value for l+w from (3) into (1)
h-1+h=17 -> h=9
Using (2), we get 2w=10 or w=5
Use (1) or (3) to get l
From (3) we have l+5=8 or l=3